If 7+5+9=18
9+8+8=8
So 9+9+9=??

Answer by srinivas.g(540)

7+5+9 =18
9+8+8=8
if we look at 7+5+9 =18
difference b/e first 2 terms =2
if you multiply 2 with third term ( 9 )we get 18 as the third term
similarly 9+8+8 = 8
difference b/e first 2 terms =1
if you multiply 1 with third term (8) we get 8 as the third term
similarly
9-9 =0 0*9 =0
hence 9+9+9=0